# Reading Between the Lines: Towards Reliable Black-box LLM Fingerprinting via Zeroth-order Gradient Estimation

### v1

@@ Abstract.
Model Fingerprinting, AI Copyright Protection, Large Language Model, Trustworthy AI
copyright: noneccs: Security and privacyccs: Computing methodologies Natural language processing
@@ 1. Introduction
Large Language Models (LLMs) have become fundamental in revolutionizing the digital landscape (Guo et al., 2025; OpenAI, 2023), powering a vast array of web applications from intelligent search engines (Sancheti et al., 2024; Xiong et al., 2024) and conversational agents (Deng et al., 2024; Ning et al., 2025) to sophisticated content creation platforms (Comanici et al., 2025). The development of these state-of-the-art (SOTA) models, however, demands substantial investment in computational resources, massive datasets, and extensive research expertise, rendering them highly valuable intellectual property for their developers (Wang and Kerschbaum, 2021; Zeng et al., 2024; Zhang et al., 2025b). Consequently, this value has given rise to significant security concerns regarding their copyright. Adversaries may engage in copyright infringement by using open-source models for commercial purposes without authorization (Shao et al., 2025a), or even resort to model stealing (Carlini et al., 2024), where proprietary models may be illicitly replicated and deployed. Protecting the copyright of these models is therefore not just a legal formality but a critical necessity to promote sustainable innovation and maintain a trustworthy AI ecosystem (Xu et al., 2025; Ren et al., 2024).
@@
To address copyright concerns, LLM fingerprinting has emerged as a promising solution (Pasquini et al., 2025; Zeng et al., 2024; Zhang et al., 2025b; Gubri et al., 2024). This technique aims to determine whether a third-party suspicious model is derived from a copyrighted source model by extracting its intrinsic and unique features ($i.e.$, a fingerprint) in a non-invasive manner (McGovern et al., 2025; Zheng et al., 2022; Chen et al., 2022). Copyright attribution is then performed by measuring the similarity between the fingerprints of the two models. Compared to invasive LLM watermarking techniques (Li et al., 2023; Xu et al., 2024; Shao et al., 2025b), which proactively embed a verifiable signal into a model’s parameters or outputs before its release, fingerprinting offers significant advantages (Shao et al., 2025a). First, it does not require model fine-tuning, which means it has a smaller overhead and also does not degrade the model’s performance. Second, it can be applied to protect models that have already been deployed or released, a scenario where watermarking is not a viable option. Consequently, LLM fingerprinting has become a more practical and mainstream solution for copyright auditing.
@@
Existing LLM fingerprinting techniques are primarily classified based on the level of access to the suspicious model, leading to two categories: white-box and black-box methods (Shao et al., 2025a). White-box fingerprinting assumes full access to a model’s internal components, including its architecture and parameters. This allows fingerprints to be extracted directly from static model weights (Zeng et al., 2024; Zheng et al., 2022), intermediate representations (Zhang et al., 2025b, 2024a), or gradients of weights (Wu et al., 2025). Conversely, black-box fingerprinting operates under the constraint that the model developer can only interact with the model through the API and observe the corresponding outputs. This paradigm is further divided into two strategies. The first, untargeted fingerprinting (Sun et al., 2025; Pasquini et al., 2025; Jin et al., 2024), involves submitting a selected set of queries to different models and comparing their outputs, where high similarity can indicate copyright infringement. The second strategy, targeted fingerprinting (Tsai et al., 2025; Gubri et al., 2024), generates a unique set of query-response pairs by optimizing them on a source model and then verifies if a suspicious model can consistently reproduce these specific outputs.
@@ 2.1. Large Language Models
Large Language Models (LLMs) have become a cornerstone of the modern web, fundamentally reshaping how users interact with online services and information (Deng et al., 2024; Zhu et al., 2025). An LLM is a sophisticated computational system, typically defined by its vast parameter scale (Guo et al., 2025; Team, 2024c; Grattafiori et al., 2024). The generation of text by these models follows an auto-regressive paradigm. This means that the model produces text sequentially, one token at a time, where each new token is conditionally dependent on the sequence of previously generated tokens. The probability of generating a sequence of tokens $Y=(y_{1},y_{2},\dots,y_{T})$ is factorized into a product of conditional probabilities:
@@
This underlying mechanism enables the diverse web applications we see today, such as web agents (Ning et al., 2025; Deng et al., 2024), search engines (Sancheti et al., 2024), and code generation (Nam et al., 2024; He et al., 2025). As LLMs become increasingly central to technological innovation and economic productivity on the web, they represent one of the most significant advancements in computing and have become immensely valuable digital assets. Consequently, protecting the intellectual property of these models within the web environment has emerged as a critical challenge.
@@ 2.2. LLM Fingerprinting
White-box LLM Fingerprinting. White-box fingerprinting methods operate under the assumption that the auditor has full access to a model’s internal components, including its architecture and parameters. These techniques are categorized based on their feature source into three different methods: static, forward-pass, and backward-pass methods (Shao et al., 2025a). Static fingerprinting directly analyzes a model’s learned parameters as the primary identifier (Zeng et al., 2024; Zheng et al., 2022). Forward-pass methods leverage the dynamic intermediate representations generated as the model processes inputs (Zhang et al., 2025b, 2024a), while backward-pass techniques use the gradients produced during backpropagation as a unique signature (Wu et al., 2025). Although effective, the practical application of these methods is limited to open-source models where model parameters have already been accessible.
@@
  • Targeted Fingerprinting: This is a more purposeful approach that seeks to construct a specific set of query-response pairs that are unique to a source model and its derivatives (Gubri et al., 2024; Tsai et al., 2025; Jin et al., 2024). The key idea is that these special pairs act as a fingerprint, and only models from the source’s lineage will be able to consistently reproduce the predefined target responses. To perform an audit, the auditor submits these optimized queries to a suspicious model and checks if its outputs match the target responses. If a significant number of responses match, the suspicious model is flagged as a likely derivative.
@@ 3.2. The Theoretical Edge: Why Gradients Carry More Information
Proof.
Please refer to Appendix A. ∎
Theorem 2 provides the mathematical proof for our hypothesis. The coefficient $c_{1}^{2}/(2W^{2}c_{2}^{2}\sigma_{X}^{2})+4>1$, leading to $\mathcal{I}_{D}(W)>\mathcal{I}_{Y}(W)$. The result formally establishes that under the given conditions, the gradient $D$ contains approximately more information about the parameter $W$ than the output $Y$ does. This theoretical advantage confirms that the gradient is an inherently more distinctive and reliable feature for a robust model fingerprint.
@@ 4.1. Threat Model
  • The auditor has only black-box, API-level access to both the source model and the suspicious model. This access is limited to providing inputs and observing the corresponding textual outputs. Internal states and output probability logits are inaccessible in our threat model.
@@ 4.4. Fingerprint Generation via Zeroth-order Gradient Estimation
where $\alpha$ is the regularization hyperparameter and $\|\cdot\|_{F}$ is the Frobenius norm. Finally, a Jacobian matrix $J_{i}$ is computed for each of the $n$ base queries. The overall fingerprint for the model $\mathcal{M}$, denoted as $J_{\mathcal{M}}$, is the element-wise average of these individual Jacobian matrices. This aggregation step creates a more comprehensive and stable fingerprint that captures the model’s characteristic behavior across a diverse set of inputs.
(11) $J_{\mathcal{M}}=\frac{1}{n}\sum_{i=1}^{n}J_{i}.$
@@ 5.1. Experimental Settings
Baselines. For the baseline methods, we implement 3 untargeted black-box methods, LLMmap (Pasquini et al., 2025), Model Equality Testing (MET) (Gao et al., 2025), and Sentence Embedding Fingerprinting (SEF) (Shao et al., 2025a), and 1 targeted black-box method, TRAP (Gubri et al., 2024). We also include a white-box method, REEF (Zhang et al., 2025b), for reference. Note that targeted methods, including TRAP, are typically semi-black-box. They assume black-box access to the suspicious model, but white-box access to the source model, which is a slightly stronger assumption than untargeted methods. The maximum number of queries is set to $q=200$.
@@ 5.2. Main Results
The results demonstrate that ZeroPrint achieves SOTA or comparable performance across all evaluated metrics. It significantly outperforms existing untargeted black-box methods like LLMmap, MET, and SEF. Notably, ZeroPrint also shows superior performance compared to the targeted, semi-black-box method TRAP, improving the AUC to 0.720 and the pAUC to 0.683. However, a performance gap remains when compared to white-box methods like REEF, which have direct access to model internals. It is worth noting that these results were achieved under a strict query limit of 200 queries per model. Subsequent experiments in our ablation study reveal that ZeroPrint’s performance can be further enhanced by increasing the number of queries allowed. This scalability is further detailed in Section 5.3.
@@ 5.3. Ablation Study
In this subsection, we study the effects of several important hyperparameters or elements in ZeroPrint.
@@
Effects of Different Query Data. The choice of data for constructing base queries is another crucial element. We evaluated the performance of ZeroPrint using base queries derived from five different datasets: HumanEval (Chen et al., 2021) (our default), three question-answering (QA) datasets (TruthfulQA (Lin et al., 2022), SQuAD (Rajpurkar et al., 2016), and DROP (Dua et al., 2019)), and one text completion dataset (Wikipedia (Foundation, [n. d.])). The results, presented in Table 2, show that the source of the query data significantly impacts fingerprinting effectiveness. When using queries from the QA datasets, the performance was consistently poor. Similarly, the Wikipedia dataset, despite being a completion task, also failed to achieve comparable results to HumanEval. This can be attributed to the nature of the tasks. Pure text completion is inherently more diverse and less structured than code completion. Thus, various derivative models tend to exhibit less consistent behavior on text-based prompts, making it harder to extract a stable and reliable fingerprint. In contrast, the structured nature of code completion elicits more uniform responses, leading to better performance.
@@ 5.4. Resistance to Potential Adaptive Attacks
Figure 3. The time overhead of different LLM fingerprinting methods.
5.5. Overhead
To evaluate practical efficiency, we measured the total time for fingerprint extraction and verification on a subset of four LLMs. Detailed Settings can be found in Appendix B. From Figure 3, ZeroPrint demonstrates a competitive overhead of 424.9 seconds, which is significantly faster than most black-box baselines like MET, SEF, and TRAP. Its runtime is only greater than the white-box REEF and LLMmap, the latter of which uses a minimal set of only 8 queries. Given that the fingerprinting for a single model can be completed in under two minutes on average, ZeroPrint’s process completes in an acceptable timeframe. Overall, ZeroPrint is a computationally practical solution for real-world auditing.
@@ Appendix A Proof of Theorem 2
(17) $\mathbb{E}\left[\left(\frac{(Z)}{\sigma^{2}(\theta)}\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\right]=$ $\frac{1}{\sigma^{4}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\mathbb{E}[(Z)^{2}]$
$=$ $\frac{1}{\sigma^{2}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}.$
@@
$=\frac{1}{W^{2}\sigma_{X}^{2}}*0^{2}+\frac{1}{2(W^{2}\sigma_{X}^{2})^{2}}(2W\sigma_{X}^{2})^{2}$
$=\frac{4W^{2}\sigma_{X}^{4}}{2W^{4}\sigma_{X}^{4}}=\frac{2}{W^{2}}.$
@@
(27) $\frac{\partial\mu_{D}(W)}{\partial W}$ $=c_{1},$
$\frac{\partial\sigma^{2}_{D}(W)}{\partial W}$ $=4W^{3}c_{2}^{2}\sigma_{X}^{2}.$
@@
$=\frac{c_{1}^{2}}{W^{4}c_{2}^{2}\sigma_{X}^{2}}+\frac{8}{W^{2}}$
$=\frac{2}{W^{2}}(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4).$
@@ Appendix B Implementation Details
1:Input:
@@
3:$n$: Number of base queries to select
4:$m$: Number of perturbed queries to generate per base query
5:$r$: Number of words to replace in each perturbation
6:$k$: Number of top semantically similar candidates to consider for replacement
[…]
8: Output:
9:$Q$: The query set containing base and perturbed queries
10:
[…]
13:for $i\leftarrow 1$ to $n$ do $\triangleright$ Select n base queries
14:  $x_{i}\leftarrow\text{SampleAndFormat}(D_{source})$ $\triangleright$ e.g., “Complete the code: [snippet]”
15:  $Q_{base}\leftarrow Q_{base}\cup\{x_{i}\}$
[…]
17:
18:for each base query $x_{i}\in Q_{base}$ do
19:  for $j\leftarrow 1$ to $m$ do $\triangleright$ Generate m perturbed versions
20:   $x^{\prime}_{i,j}\leftarrow x_{i}$ $\triangleright$ Create a copy of the base query
21:   $W_{replaceable}\leftarrow\text{SelectRandomWords}(x_{i},r)$ $\triangleright$ Select r words to replace
22:   for each word $w\in W_{replaceable}$ do
23:     $C_{candidates}\leftarrow E_{word}.\text{TopKSimilar}(w,k)$ $\triangleright$ Find top-k similar words
24:     $w_{replacement}\leftarrow\text{RandomChoice}(C_{candidates})$
25:     $x^{\prime}_{i,j}\leftarrow\text{ReplaceWord}(x^{\prime}_{i,j},w,w_{replacement})$
26:   $Q\leftarrow Q\cup\{x^{\prime}_{i,j}\}$ $\triangleright$ Add the perturbed query to the final set
27:return $Q$
Algorithm 2 ZeroPrint Fingerprint Generation
1:Input:
[…]
6:$t$: Number of times to query the model for each input to average outputs
[…]
8: Output:
[…]
10:
[…]
12:for each base query $x_{i}\in Q_{base}$ do
13:  $\Delta X_{i},\Delta Y_{i}\leftarrow[],[]$ $\triangleright$ Initialize lists for input/output differences
14:  $\triangleright$ Get stable embedding for the base query’s output
15:  $\{y_{i,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x_{i},t)$
16:  $\bar{e}_{y_{i}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E_{sent}(y_{i,k})$
17:  $e_{x_{i}}\leftarrow E_{sent}(x_{i})$
18:  for each perturbed query $x^{\prime}_{i,j}$ for $x_{i}$ do
19:   $\triangleright$ Get stable embedding for the perturbed query’s output
20:   $\{y^{\prime}_{i,j,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x^{\prime}_{i,j},t)$
21:   $\bar{e}_{y^{\prime}_{i,j}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E_{sent}(y^{\prime}_{i,j,k})$
22:   $e_{x^{\prime}_{i,j}}\leftarrow E_{sent}(x^{\prime}_{i,j})$
23:   $\triangleright$ Calculate input and output embedding differences
24:   $\Delta x_{i,j}\leftarrow e_{x^{\prime}_{i,j}}-e_{x_{i}}$
25:   $\Delta y_{i,j}\leftarrow\bar{e}_{y^{\prime}_{i,j}}-\bar{e}_{y_{i}}$
26:   Append $\Delta x_{i,j}$ to $\Delta X_{i}$
27:   Append $\Delta y_{i,j}$ to $\Delta Y_{i}$
28:  $\triangleright$ Estimate the local Jacobian using Ridge Regression
[…]
30:  Append $J_{i}$ to $J_{list}$
[…]
33:return $J_{M}$
Detailed Settings for the Baseline Fingerprinting Methods. In our experiments, we take 1 white-box method and 4 state-of-the-art black-box methods as our baseline. For fair comparison, following (Zhang et al., 2025b), we also set the query limit to $q=200$ while applying these methods. The detailed settings are as follows.
  • REEF (Zhang et al., 2025b): REEF is a forward-pass white-box fingerprinting method that relies on the intermediate representations as fingerprints. We strictly follow the original setting and sample 200 queries from TruthfulQA (Lin et al., 2022) as the input queries.

### v2

@@ Abstract.
Model Fingerprinting; AI Copyright Protection; Large Language Model; Trustworthy AI
journalyear: 2026conference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792196isbn: 979-8-4007-2307-0/2026/04ccs: Security and privacyccs: Computing methodologies Natural language processing
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18325480 and https://github.com/shaoshuo-ss/ZeroPrint.
@@ 1. Introduction
Large Language Models (LLMs) have become fundamental in revolutionizing the digital landscape (OpenAI, 2023; Guo et al., 2025), powering a vast array of web applications from intelligent search engines (Sancheti et al., 2024; Xiong et al., 2024) and conversational agents (Deng et al., 2024; Ning et al., 2025) to sophisticated content creation platforms (Comanici et al., 2025). However, the development of these state-of-the-art (SOTA) models demands substantial investment in computational resources, massive datasets, and extensive research expertise, rendering them highly valuable intellectual property for their developers (Wang and Kerschbaum, 2021; Zeng et al., 2024; Zhang et al., 2025a). Consequently, this value has given rise to significant security concerns regarding their copyright. Adversaries may engage in copyright infringement by using open-source models for commercial purposes without authorization (Shao et al., 2025a), or even resort to model stealing (Carlini et al., 2024; Yao et al., 2025), where proprietary models may be illicitly replicated and deployed. Protecting the copyright of these models is therefore not just a legal formality but a critical necessity to promote sustainable innovation and maintain a trustworthy AI ecosystem (Xu et al., 2025; Ren et al., 2024).
@@
To address copyright concerns, LLM fingerprinting has emerged as a promising solution (Pasquini et al., 2025; Zeng et al., 2024; Zhang et al., 2025a; Gubri et al., 2024). This technique aims to determine whether a third-party suspicious model is derived from a copyrighted source model by extracting its intrinsic and unique features ($i.e.$, a fingerprint) in a non-invasive manner (McGovern et al., 2025; Zheng et al., 2022; Chen et al., 2022). Copyright attribution is then performed by measuring the similarity between the fingerprints of the two models. Compared to invasive LLM watermarking techniques (Li et al., 2023; Xu et al., 2024; Shao et al., 2025b), which proactively embed a verifiable signal into a model’s parameters or outputs before its release, fingerprinting offers significant advantages (Shao et al., 2025a). First, it does not require model fine-tuning, which means it has a smaller overhead and also does not degrade the model’s performance. Second, it can be applied to protect models that have already been deployed or released, a scenario where watermarking is not a viable option. Consequently, LLM fingerprinting has become a more practical and mainstream solution for copyright auditing.
@@
Existing LLM fingerprinting techniques are primarily classified based on the level of access to the suspicious model, leading to two categories: white-box and black-box methods (Shao et al., 2025a). White-box fingerprinting assumes full access to a model’s internal components, including its architecture and parameters. This allows fingerprints to be extracted directly from static model weights (Zeng et al., 2024; Zheng et al., 2022), intermediate representations (Zhang et al., 2025a, 2024a), or gradients of weights (Wu et al., 2025). Conversely, black-box fingerprinting operates under the constraint that the model developer can only interact with the model through the API and observe the corresponding outputs. This paradigm is further divided into two strategies. The first, untargeted fingerprinting (Sun et al., 2025; Pasquini et al., 2025; Jin et al., 2024), involves submitting a selected set of queries to different models and comparing their outputs, where high similarity can indicate copyright infringement. The second strategy, targeted fingerprinting (Tsai et al., 2025; Gubri et al., 2024), generates a unique set of query-response pairs by optimizing them on a source model and then verifies if a suspicious model can consistently reproduce these specific outputs.
@@ 2.1. Large Language Models
Large Language Models (LLMs) have become a cornerstone of the modern web, fundamentally reshaping how users interact with online services and information (Deng et al., 2024; Zhu et al., 2025; Luo et al., 2026). An LLM is a sophisticated computational system, typically defined by its vast parameter scale (Guo et al., 2025; Team, 2024c; Grattafiori et al., 2024). The generation of text by these models follows an auto-regressive paradigm. This means that the model produces text sequentially, one token at a time, where each new token is conditionally dependent on the sequence of previously generated tokens. The probability of generating a sequence of tokens $Y=(y_{1},y_{2},\dots,y_{T})$ is factorized into a product of conditional probabilities:
@@
This underlying mechanism enables the diverse web applications, such as search engines (Sancheti et al., 2024), code generation (Nam et al., 2024; He et al., 2025), and web agents (Ning et al., 2025; Deng et al., 2024). LLMs represent one of the most significant advancements in computing and have become immensely valuable digital assets. Consequently, protecting the copyright of these models within the web environment has emerged as a critical challenge.
@@ 2.2. LLM Fingerprinting
White-box LLM Fingerprinting. White-box fingerprinting methods operate under the assumption that the auditor has full access to a model’s internal components, including its architecture and parameters. These techniques are categorized based on their feature source into three different methods: static, forward-pass, and backward-pass methods (Shao et al., 2025a). Static fingerprinting directly analyzes a model’s learned parameters as the primary identifier (Zeng et al., 2024; Zheng et al., 2022). Forward-pass methods leverage the dynamic intermediate representations generated as the model processes inputs (Zhang et al., 2025a, 2024a), while backward-pass techniques use the gradients produced during backpropagation as a unique signature (Wu et al., 2025). Although effective, the practical application of these methods is limited to open-source models where model parameters have already been accessible.
@@
  • Targeted Fingerprinting: This paradigm seeks to construct a specific set of query-response pairs that are unique to a source model and its derivatives (Gubri et al., 2024; Tsai et al., 2025; Jin et al., 2024). The key idea is that these special pairs act as a fingerprint, and only models from the source’s lineage will be able to consistently reproduce the predefined target responses. To perform an audit, the auditor submits these optimized queries to a suspicious model and checks if its outputs match the target responses. If a significant number of responses match, the model will be flagged as a likely derivative.
@@ 3.2. The Theoretical Edge: Why Gradients Carry More Information
The proof is in Appendix A. Theorem 2 provides the mathematical proof for our hypothesis. The coefficient $c_{1}^{2}/(2W^{2}c_{2}^{2}\sigma_{X}^{2})+4>1$, leading to $\mathcal{I}_{D}(W)>\mathcal{I}_{Y}(W)$. The result formally establishes that the gradient $D$ contains approximately more information about the parameter $W$ than the output $Y$ does. This theoretical advantage confirms that the gradient is an inherently more distinctive and reliable feature for a robust model fingerprint.
@@ 4.1. Threat Model
  • The auditor has only black-box, API-level access to both the source and suspicious models. This access is limited to providing inputs and observing the corresponding textual outputs. Internal states and output logits are inaccessible in our threat model.
@@ 4.4. Fingerprint Generation via Zeroth-order Gradient Estimation
where $\alpha$ is the regularization hyperparameter and $\|\cdot\|_{F}$ is the Frobenius norm. Finally, a Jacobian matrix $J_{i}$ is computed for each of the $n$ base queries. The overall fingerprint for the model $\mathcal{M}$, denoted as $J_{\mathcal{M}}$, is the element-wise average of these individual Jacobian matrices. This aggregation step creates a more comprehensive and stable fingerprint that captures the model’s characteristic behavior across a diverse set of inputs: $J_{\mathcal{M}}=\frac{1}{n}\sum_{i=1}^{n}J_{i}$.
@@ 5.1. Experimental Settings
Baselines. For the baseline methods, we implement 3 untargeted black-box methods, LLMmap (Pasquini et al., 2025), Model Equality Testing (MET) (Gao et al., 2025), and Sentence Embedding Fingerprinting (SEF) (Shao et al., 2025a), and 1 targeted black-box method, TRAP (Gubri et al., 2024). We also include a white-box method, REEF (Zhang et al., 2025a), for reference. Note that targeted methods, including TRAP, are typically semi-black-box. They assume black-box access to the suspicious model, but white-box access to the source model, which is a slightly stronger assumption than untargeted methods. The maximum number of queries is set to $q=200$.
@@ 5.2. Main Results
The results demonstrate that ZeroPrint achieves SOTA or comparable performance across all evaluated metrics. It significantly outperforms existing untargeted methods. Notably, ZeroPrint also shows superior performance compared to the targeted, semi-black-box method TRAP, improving the AUC to 0.720 and the pAUC to 0.683. However, a performance gap remains when compared to white-box methods like REEF, which have direct access to model internals. It is worth noting that these results were achieved under a strict query limit of 200 queries per model. Subsequent experiments in our ablation study reveal that ZeroPrint’s performance can be further enhanced by increasing the number of queries allowed. This scalability is further detailed in Section 5.3.
@@ 5.3. Ablation Study
Effects of Different Query Data. The choice of data for constructing base queries is another crucial element. We evaluated the performance of ZeroPrint using base queries derived from five different datasets: HumanEval (Chen et al., 2021) (our default), three question-answering (QA) datasets (TruthfulQA (Lin et al., 2022), SQuAD (Rajpurkar et al., 2016), and DROP (Dua et al., 2019)), and one text completion dataset (Wikipedia (Foundation, )). The results, presented in Table 2, show that the source of the query data significantly impacts fingerprinting effectiveness. When using queries from the QA datasets, the performance was consistently poor. Similarly, the Wikipedia dataset, despite being a completion task, also failed to achieve comparable results to HumanEval. This can be attributed to the nature of the tasks. Pure text completion is inherently more diverse and less structured than code completion. Thus, various derivative models tend to exhibit less consistent behavior on text-based prompts, making it harder to extract a stable and reliable fingerprint. In contrast, the structured nature of code completion elicits more uniform responses, leading to better performance.
@@ 5.4. Resistance to Potential Adaptive Attacks
Figure 3. The overhead of different fingerprinting methods.
5.5. Overhead Evaluation
To evaluate practical efficiency, we measured the total time for fingerprint extraction and verification on a subset of four LLMs. Experimental details can be found in Appendix B. From Figure 3, ZeroPrint demonstrates a competitive overhead of 424.9 seconds, which is significantly faster than most black-box baselines like MET, SEF, and TRAP. Its runtime is only greater than the white-box REEF and LLMmap, the latter of which uses a minimal set of only 8 queries. Given that the fingerprinting for a single model can be completed in under two minutes on average, ZeroPrint’s process completes in an acceptable timeframe. Overall, ZeroPrint is a computationally practical solution for real-world auditing.
@@ 6. Conclusion
Acknowledgments
This research is supported in part by the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01169), the Kunpeng-Ascend Science and Education Innovation Excellence/Incubation Center, and the National Natural Science Foundation of China under Grants (62441238, U2441240).
@@ Appendix A Proof of Theorem 2
(16) $\mathbb{E}\left[\left(\frac{(Z)}{\sigma^{2}(\theta)}\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\right]$
$=$ $\frac{1}{\sigma^{4}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}\mathbb{E}[(Z)^{2}]=\frac{1}{\sigma^{2}(\theta)}\left(\frac{\partial\mu(\theta)}{\partial\theta}\right)^{2}.$
@@
$=$ $\frac{1}{W^{2}\sigma_{X}^{2}}*0^{2}+\frac{1}{2(W^{2}\sigma_{X}^{2})^{2}}(2W\sigma_{X}^{2})^{2}=\frac{4W^{2}\sigma_{X}^{4}}{2W^{4}\sigma_{X}^{4}}=\frac{2}{W^{2}}.$
@@
(26) $\frac{\partial\mu_{D}(W)}{\partial W}$ $=c_{1},\ \ \frac{\partial\sigma^{2}_{D}(W)}{\partial W}=4W^{3}c_{2}^{2}\sigma_{X}^{2}.$
@@
$=\frac{c_{1}^{2}}{W^{4}c_{2}^{2}\sigma_{X}^{2}}+\frac{8}{W^{2}}=\frac{2}{W^{2}}(\frac{c_{1}^{2}}{2W^{2}c_{2}^{2}\sigma_{X}^{2}}+4).$
@@
0:  $D_{source}$: Source dataset for base queries ($e.g.$, HumanEval)   $n$: Number of base queries to select   $m$: Number of perturbed queries to generate per base query   $r$: Number of words to replace in each perturbation   $k$: Number of top semantically similar candidates to consider   $E_{word}$: Pre-trained word embedding model ($e.g.$, GloVe)
0:  $Q$: The query set containing base and perturbed queries
1:  $Q_{b}\leftarrow\emptyset$ # Initialize set for base queries
2:  $Q\leftarrow\emptyset$ # Initialize the final query set
3:  for $i\leftarrow 1$ to $n$ do
4:   $x_{i}\leftarrow\text{SampleAndFormat}(D_{source})$ # e.g., “Complete the code: [snippet]”
5:   $Q_{b}\leftarrow Q_{b}\cup\{x_{i}\}$
6:  end for
7:  $Q\leftarrow Q\cup Q_{b}$ # Add all base queries to the final set
8:  for each base query $x_{i}\in Q_{b}$ do
9:   for $j\leftarrow 1$ to $m$ do
10:    $x^{\prime}_{i,j}\leftarrow x_{i}$ # Create a copy of the base query
11:    $W_{replaceable}\leftarrow\text{SelectRandomWords}(x_{i},r)$ # Select r words to replace
12:    for each word $w\in W_{replaceable}$ do
13:     $C_{candidates}\leftarrow E_{word}.\text{TopKSimilar}(w,k)$ # Find top-k similar words
14:     $w_{replacement}\leftarrow\text{RandomChoice}(C_{candidates})$
15:     $x^{\prime}_{i,j}\leftarrow\text{ReplaceWord}(x^{\prime}_{i,j},w,w_{replacement})$
16:    end for
17:    $Q\leftarrow Q\cup\{x^{\prime}_{i,j}\}$ # Add the perturbed query to the final set
18:   end for
19:  end for
20:  return $Q$
Algorithm 2 ZeroPrint Fingerprint Generation
0:  $\mathcal{M}$: The target black-box LLM   $Q_{b}=\{x_{i}\}_{i=1}^{n}$: Set of base queries   $Q_{p}=\{\{x^{\prime}_{i,j}\}_{j=1}^{m}\}_{i=1}^{n}$: Set of perturbed queries   $E$: Pre-trained sentence embedding model   $t$: Number of times to query the model for each input   $\alpha$: Regularization hyperparameter for Ridge Regression
0:  $J_{\mathcal{M}}$: The final fingerprint (aggregated Jacobian matrix)
1:  $J_{list}\leftarrow[]$ # Initialize list to store local Jacobians
2:  for each base query $x_{i}\in Q_{b}$ do
3:   $\Delta X_{i},\Delta Y_{i}\leftarrow[],[]$ # Initialize lists for input/output differences
4:   # Get stable embedding for the base query’s output
5:   $\{y_{i,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x_{i},t)$
6:   $\bar{e}_{y_{i}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E(y_{i,k})$
7:   $e_{x_{i}}\leftarrow E(x_{i})$
8:   for each perturbed query $x^{\prime}_{i,j}$ for $x_{i}$ do
9:    # Get stable embedding for the perturbed query’s output
10:    $\{y^{\prime}_{i,j,k}\}_{k=1}^{t}\leftarrow\text{QueryModelMultipleTimes}(\mathcal{M},x^{\prime}_{i,j},t)$
11:    $\bar{e}_{y^{\prime}_{i,j}}\leftarrow\frac{1}{t}\sum_{k=1}^{t}E(y^{\prime}_{i,j,k})$
12:    $e_{x^{\prime}_{i,j}}\leftarrow E(x^{\prime}_{i,j})$
13:    # Calculate input and output embedding differences
14:    $\Delta x_{i,j}\leftarrow e_{x^{\prime}_{i,j}}-e_{x_{i}}$
15:    $\Delta y_{i,j}\leftarrow\bar{e}_{y^{\prime}_{i,j}}-\bar{e}_{y_{i}}$
16:    Append $\Delta x_{i,j}$ to $\Delta X_{i}$
17:    Append $\Delta y_{i,j}$ to $\Delta Y_{i}$
18:   end for
19:   # Estimate the local Jacobian using Ridge Regression
@@
21:   Append $J_{i}$ to $J_{list}$
22:  end for
23:  $J_{\mathcal{M}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}J_{list}[i]$
24:  return $J_{\mathcal{M}}$
Appendix B Implementation Details
Detailed Settings for the Baseline Fingerprinting Methods. In our experiments, we take 1 white-box method and 4 state-of-the-art black-box methods as our baseline. For fair comparison, following (Zhang et al., 2025a), we also set the query limit to $q=200$ while applying these methods. The detailed settings are as follows.
  • REEF (Zhang et al., 2025a): REEF is a forward-pass white-box fingerprinting method that relies on the intermediate representations as fingerprints. We strictly follow the original setting and sample 200 queries from TruthfulQA (Lin et al., 2022) as the input queries.